OpenStudy (rational):

If the length of one leg of a right triangle is 16 and the length of the other leg is odd, what are the lengths of all the sides of the triangle?

OpenStudy (perl):

thats implied by 'odd' :)

OpenStudy (perl):

ill brute force it , but theres probably an elegant solution

OpenStudy (perl):

the sides are 16, 63 and

OpenStudy (perl):

65

OpenStudy (rational):

brute force isnt bad i guess as the numbers are small.. yes its giving 63, 65 haha

OpenStudy (rational):

it was asked by another user sometime ago..i was stuck at finding the solution because i didnt try brute force but yeah im also pretty sure ther exists some analytic solution

OpenStudy (perl):

There is a general formula for integer solutions to a^2 + b^2 = c^2

OpenStudy (perl):

the 'primitive' pythagorean triples, a,b are not both even (you can factor out an even)

OpenStudy (rational):

I see... a, b must have opposite parity. even-odd or odd-even

OpenStudy (rational):

(when a,b,c are primitive)

OpenStudy (perl):

i can look in my book, theres a chapter on it, somewhere

OpenStudy (rational):

no its okay.. im just messing wid this problem as im free right now and this looks interesting..

OpenStudy (perl):

my brute force was pretty lame. i used a TI 84 and did sqrt( (2x+1)^2 + 16^2), then i looked on the column for integers

OpenStudy (perl):

looked on the y column, scrolled down

OpenStudy (kainui):

Since even + odd = odd I figured out this formula: \[\Large 2^8+(2n+1)^2=(2m+1)^2 \\ \Large 2^6 +n^2+n=m^2+m \\ \Large 2^6 = (m+n+1)(m-n) \\ \Large 2^a=(m+n+1) \text{ and } 2^b=(m-n) \\ \Large a+b = 6 \text{ and } a>b\] However I noticed each of these are off in their parity. So b=0 must be true. So adding and subtracting the two equations to solve for m and n yields \[\Large 2^6=m+n-1 \\ \Large 1=m-n \\ \Large 2^5 = m \\ \Large 2^5-1=n\] Now it's plug in time. =) \[\Large 2^6-1= 63\\ \Large 2^6+1 = 65\]

OpenStudy (kainui):

I found an even better solution!\[\Large 2^8+b^2=c^2 \\ \Large 2^8 = (c+b)(c-b) \\ \Large 2^x = c+b \text{ and } 2^y = c-b \text{ with } x+y=8 \\ \Large c=\frac{2^x+2^y}{2} \text{ and } b=\frac{2^x-2^y}{2} \] Now we know y=1 since b and c are both odd. \[\Large c= 2^6+1 \\ \Large b=2^6-1\]

OpenStudy (rational):

Nice :)

OpenStudy (kainui):

This was a good question, haha got a lot of fun out of it.

OpenStudy (rational):

me too i tried generating formula and missed most of above fun haha

OpenStudy (kainui):

Oh I'd be curious to see what you tried if you want to share.

OpenStudy (rational):

a^2 + b^2 = c^2 2xy = 16 = b x^2-y^2 = a x^2 + y^2 = c ------------------------------- from first equation we have xy=8, that implies x = 8 and y = 1 therefore a = x^2 - y^2 = 8^2 - 1 = 63 c = x^2 + y^2 = 8^2 + 1 = 65

OpenStudy (rational):

that relies on identity (x^2-y^2)^2 + (2xy)^2 = (x^2+y^2)^2

OpenStudy (kainui):

Oh ok so is that the easiest form to remember for generating pythagorean triples? I know there exists a formula or system of equations or something to find them but I always have to look it up.

OpenStudy (rational):

above formula generates all primitive pythagorean triples i think proof shouldnt be hard..

OpenStudy (rational):

we may try proof by contradiction... and i feel thats as simple/elementary as it can possibly get... yeah there are many other interesting ways to generate PTs though

OpenStudy (mathteacher1729):

@Kainui those are nice solutions! :)